Shift summary for eve, Aug. 27, 1999

Angela Bellavance (bellavan@rice.edu)
Sat, 28 Aug 1999 00:06:51 -0500 (CDT)

Shift Summary for eve shift, 27/Aug/99

On Shift: Theo Alexopoulos, Angela Bellavance

Hours of beam delivered at all intensities: 8.0/8.5
Hours of beam delivered at nominal intensity: 7.9/8.5
Hours of E832 data (special runs not included): 4.8/8.5
Hours of special runs (laser scans, muon runs...): 0.0/8.5

Accounting of lost time: trigger latch crate 8 problems, TeV quench

========================================================================
Runs ended this shift:
Run 14346 - 6.4M events - BEAM_832
Run 14347 thru 14350 - crap, not written out
Run 14351 - in progress - BEAM_8332

Tapes brought to FCC this shift: QKC333,QKC334,QKC335,
QKK333,QKK334,QKK335,
QKS333,QKS334,QKS335,
QKB116
========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
Smooth data taking after major problems at the beginning.

Explaination by RickK of what happened with the Ktevana errors at the
beginning of this shift:

Trigger latch cr 8 [deep latches] gave incessant ktevana errors at
the end of the day shift. We eventaully traced it to a bad CTI;
the CTI was swapped and it seemed to work, at least with only DYC
and CTI in the readout. When the four deep latches were put back in
the readout, the PERMIT was always held in that crate. We finally
gave up and decided to bypass it and hope for the best until the next
access. We lose some trigger monitoring for the weekend.

We began the next run with tons of CsI problems, and then realized
that the latched SPILL signal was gone, so the csi monitoring thought
that all on-spill pedestals were off-spill. This is bad. So we ran
a spill signal into a spare AZLAT input in crate 53 (shared with
user word and REG pos signals), and then tweeked Vals CHKONSL subroutine
to check the AZLAT spill as well. Miraculously it worked and Val's
spirit is happy and so are we. Elliott will modify ktevana so that
the AZLAT spill bit gets copied to the missing deep latch bit and
you will never need to know that any of this horror actually happened.

Runs affected begin with 14351. Note that many monitor plots will
look bad, such as DCOR efficiency of ZERO! Please use some judgement
before paging people on OWL to tell them that their plots look bad.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14346 145 15:30 Take over shift with Ktevana errors all over the
place.
15:40 Decide to stop run and change tapes while errors
are being investigated by RickK. See note in
"Overview" of what happened with this error.
15:50 Tape mount request granted: QKNL27
17:00 We notice the humidity is high downstairs and in
the counting room. Page Leonard.
17:30 No response from Leonard. Call MCR who sends a
maintenance person.
17:45 Maintenance person arrives, says he can't find
anything tripped and the compressors all seem to
be working. He suggests we just keep trying to
contact Leonard.
14351 1 18:30 We are finally taking data again!
23 19:02 Ktevana errors - KFREND.
40 19:25 Valeri sends some kind of TRD signal that makes
the triggers hang. Did a kfrend_dyc_cti_lat and
everything's working again.
19:50 Three guys from Stanford who work on MINOS stopped
by to visit Theo. He gives them a tour but
forgets to have them sign the guest book.
66 20:01 Beam abort and beam thereafter disappears.
Channel 13 reports TeV quench at F23, recovering.
70 20:33 Beam returns! Targeting is off for two spills.
148 21:51 Beam gone for one spill.
22:30 Gas checklist done. Shift checklist done.
204 23:05 Ktevana errors - KFREND.
232- 23:43 Trigger hangs. Crate 15 is full. Quick pipeline
240 reset doesn't work. Full pipeline w/DBC plus a
couple KFRENDs works.
243 23:58 Channel #450 has gain drift of -15. Paged Sydney
and turned over to BobT who is on the owl shift.
00:00 yield832 => 3.47e6

========================================================================
Follow up Notes on Previous Shift

Did a drt_fresh_start and display zap but it didn't help much... :)

========================================================================
Items to Follow up on Next Shift

No muon run needed tonight!

Again, some of the online plots will look weird with the new trigger
configuration. They are marked in the plot book with sticky-notes.

For Theo,
_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke